OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form.

OpenStudy (anonymous):

A. 4x + 4y = 6 B. x + 4y = 12 C. 4x + 4y = -12 D. x - 4y = -6

OpenStudy (anonymous):

@FireKat97

OpenStudy (firekat97):

Okay so this time, the only thing thats changed is the fact that you have no longer been given the slope, but remember you can find the slope with m = (y2 - y1)/(x2 - x1) or just rise/run

OpenStudy (firekat97):

And from there you can apply the same equation you had used earlier

OpenStudy (anonymous):

uhh..

OpenStudy (carlyleukhardt):

Let (2, 2) and (6, 3) be (x1,y1) & (x2,y2) First find the slope m=(y2-y1)/(x2-x1) m=1/4 Then take any pt (2, 2) or (6, 3) (y-y1)=m(x-x1) Then substitute the values of m & x1&y1 x-4y+6=0

OpenStudy (firekat97):

okay have you covered the equation m = (y2 - y1)/(x2 - x1)?

OpenStudy (anonymous):

okay so 3-2=1 and 6-2=4

OpenStudy (carlyleukhardt):

yes

OpenStudy (firekat97):

yup but its (3 - 2)/(6 -2) so m becomes 1/4

OpenStudy (anonymous):

indeed whats next

OpenStudy (firekat97):

and now we have our slope and you can pick either of the two set of coordiantes, and reapply the formula we used in the previous question

OpenStudy (firekat97):

yup so in this case (2, 2) looks like the easier coordinate to use (because smaller numbers) so you can sub this and the slop of 1/4 into (y - y1) = m(x - x1)

OpenStudy (firekat97):

slope*

OpenStudy (anonymous):

(y-2) =1/4 (x-2)

OpenStudy (firekat97):

yup thats correct!

OpenStudy (anonymous):

thats not a choice though..

OpenStudy (firekat97):

Thats because they don't have fractions in the answer, so we should rearrange our equation

OpenStudy (firekat97):

so to get rid of the fraction on the right hand side, we can multiply both sides of the equation by 4 for starters

OpenStudy (anonymous):

4y-8=x-2

OpenStudy (firekat97):

Yup, and now we simplify further by resolving the -8 on the left hand side and -2 on the right

OpenStudy (anonymous):

4y=x+6

OpenStudy (anonymous):

4y-x=6

OpenStudy (firekat97):

yup thats correct!

OpenStudy (anonymous):

thats not an answer either :/

OpenStudy (firekat97):

Thats because they have multiplied the left and right hand sides by -1, so try that and tell me what you get

OpenStudy (anonymous):

d

OpenStudy (firekat97):

Yup

OpenStudy (firekat97):

:D

OpenStudy (anonymous):

Thank you!!!

OpenStudy (firekat97):

No problem :)